Jessica is 21 years older than Gabriela. Nine years ago, Jessica was 4 times as old as Gabriela.
How old is jessica now?

Answer by TimothyLamb(4379)

x = jess age now
y = gaby age now
---
x = y + 21
x - 9 = 4(y - 9)
---
x = y + 21
x - 9 = 4y - 36
---
put the system of linear equations into standard form
---
x - y = 21
x - 4y = -27
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = jess age now = 37
y = gaby age now = 16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php